OpenStudy (epicteatime):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (- 3, - 25), and x intercepts at x = -8 and x = 2. (Do not include the negative sign in your answer.)

OpenStudy (epicteatime):

@matricked

OpenStudy (anonymous):

assume y = ax^2+bx+c given y=-25 when x=-3 and y=0 when x= -8 y=0 when x=2

OpenStudy (epicteatime):

mmhhmmm?

OpenStudy (anonymous):

Alternately y =k (x+8)(x-2) you just need to find k using (- 3, - 25),

OpenStudy (epicteatime):

how do i do that?

OpenStudy (imstuck):

Find k like this, using (-3, -25)...

OpenStudy (epicteatime):

yes?

OpenStudy (epicteatime):

i want all the help i can get ;L

OpenStudy (imstuck):

In y = k(x+8)(x-2), your x values will be -3, and your y value will be -25 (from your point (-3, -25)) Fill those in and solve for k:

OpenStudy (epicteatime):

i got -25=k(5-5) ;P

OpenStudy (imstuck):

-25=k(-3+8)(-3-2) -25=k(5)(-5) -25=-25k k=1

OpenStudy (epicteatime):

oh

OpenStudy (epicteatime):

so its...

OpenStudy (imstuck):

You had the right idea above, you just put the two 5s in the same set of parenthesis instead of in their own andd multiplying them.

OpenStudy (epicteatime):

so its -1?

OpenStudy (imstuck):

When you find k, or a as I like to call it, you fill it into your standard equation\[ax ^{2}+bx+c\]like this:

OpenStudy (epicteatime):

yes?

OpenStudy (imstuck):

Use the roots as your factors. We have roots of -8 and 2. So the roots are (x + 8)(x - 2), right? Do this, then, with the k (or a) value:

OpenStudy (epicteatime):

ok

OpenStudy (imstuck):

a(x+8)(x-2)-->1((x+8)(x-2)-->x^2+6x-16

OpenStudy (epicteatime):

yes

OpenStudy (imstuck):

That's your equation in standard form...\[y=x ^{2}+6x-16\]

OpenStudy (epicteatime):

ohhh

OpenStudy (epicteatime):

thx man ;3

OpenStudy (imstuck):

The clue is, find your "a" or "k" (same thing) by using your vertex and your roots. Then use the roots and multiply them by your a value to expand.

OpenStudy (epicteatime):

now if you could help me with a quadratic equation i made, and tell me if you can solve it using the factorisation, that would be really helpful ;3

OpenStudy (epicteatime):

ohh

OpenStudy (imstuck):

And BTW...I'm not a man, I'm a ma'am. But that's ok!

OpenStudy (epicteatime):

2a^2+5x+14

OpenStudy (epicteatime):

oh sry ma'am

OpenStudy (epicteatime):

anyways that was the equation

OpenStudy (imstuck):

you can't factor that because you have both an a and an x in it. Which is it? x or a?

OpenStudy (epicteatime):

oh sry

OpenStudy (epicteatime):

2x^2+5x+14

OpenStudy (epicteatime):

is that factorable?

OpenStudy (imstuck):

With the quadratic formula, anything is factorable!!!

OpenStudy (epicteatime):

i need an equation that can't be factorised ;P

OpenStudy (epicteatime):

care to give me an example?

OpenStudy (imstuck):

Ok, then, that can't be factored. The solutions when you use the quadratic formula are\[\frac{ -5\pm i \sqrt{87} }{ 4 }\]That is a polynomial with nonreal complex solutions. I'd say it would be fair to say that doesn't factor. Not at least the "regular" way! You can use it!

OpenStudy (epicteatime):

yay!

OpenStudy (epicteatime):

thank you!

OpenStudy (imstuck):

Yay is right!

OpenStudy (epicteatime):

huzzah!

OpenStudy (imstuck):

You're welcome!